### v1

@@ 2.3 Exponential and logarithmic mapping
Ganea et al. [38] formalized hybrid operations for hyperbolic neural networks using logarithmic and exponential mappings, establishing a principled framework for bridging Euclidean and hyperbolic computations. This approach enables the adaptation of standard neural network operations to hyperbolic manifolds through a systematic mapping strategy. Specifically, it leverages the tangent space at the origin of the Poincaré ball, which is isometric to Euclidean space $\mathbb{R}^{n}$. The logarithmic map projects points from the hyperbolic manifold to this tangent space by
(5) $\text{log}_{0}(\mathbf{x})=\text{artanh}(\|\mathbf{x}\|)\frac{\mathbf{x}}{\|\mathbf{x}\|},$
with the inverse exponential map returning points to the hyperbolic space using
(6) $\text{exp}_{0}(\mathbf{v})=\tanh(\|\mathbf{v}\|)\frac{\mathbf{v}}{\|\mathbf{v}\|}.$
These mappings preserve distances in the radial direction and enable a workflow in which (i) hyperbolic points are mapped to the tangent space via $\text{log}_{0}$, (ii) standard Euclidean operations are performed, and (iii) results are mapped back to hyperbolic space via $\text{exp}_{0}$. To enable direct computation in hyperbolic space, Ungar [104, 103] introduced Möbius gyrovector algebra, which Ganea et al. [39] highlighted and utilized to generalize vector space operations to the Poincaré ball. The key operations include:
Möbius addition: For $\mathbf{x},\mathbf{y}\in\mathcal{B}_{r}^{n}$,
Möbius scalar multiplication:
These operations preserve the hyperbolic structure while enabling neural network computations such as linear transformations and activations.
@@ 2.4 Optimization techniques in Riemannian manifold
Training large models in hyperbolic space requires specialized optimization techniques, since standard Euclidean gradient descent does not directly apply to curved manifolds. Notably, numerical instabilities such as overflow, underflow, and NaNs can arise in both the Poincaré ball and Lorentz models when points approach the boundary or stray far from the origin. To combat these issues, researchers have proposed reparameterization tricks – for example, performing updates in Euclidean parameter space and mapping them into hyperbolic space – to keep representations numerically well-behaved. Furthermore, optimization algorithms have been extended to the hyperbolic realm: Riemannian SGD (RSGD) and Riemannian Adam (RADAM) update weights in the tangent space at each point on the manifold and then use the exponential map to retract back to hyperbolic space [38]. Ganea et al.[39] introduced the Möbius gyrovector framework (e.g., Möbius addition and scalar multiplication) to define neural network operations entirely in the Poincaré ball model, enabling the construction and training of hyperbolic feed-forward and recurrent networks. A key insight is to perform gradient updates intrinsically on the manifold via Riemannian optimization [68].
Riemannian Stochastic Gradient Descent (RSGD): Riemannian Stochastic Gradient Descent (RSGD) generalizes SGD to Riemannian manifolds by using exponential mappings to update parameters while maintaining manifold constraints [9]:
@@
Riemannian Stochastic Variance Reduced Gradient (RSVRG): Building on this foundation, Zhang et al. [127] introduced Riemannian SVRG (RSVRG), the first variance-reduced stochastic optimization method for Riemannian manifolds. It applies the “Stochastic Variance Reduced Gradient (SVRG)” technique in non-Euclidean settings to accelerate convergence. RSVRG combines geometric insights with variance reduction techniques, achieving linear convergence for geodesically convex functions and $\mathcal{O}(1/T)$ convergence for non-convex objectives [127, 68].
Riemannian Adagrad (RADAGRAD): Recently, RADAGRAD [43] extended adaptive gradient methods to Riemannian manifolds. For a product manifold $\mathcal{M}=\mathcal{M}_{1}\times\cdots\times\mathcal{M}_{n}$, the parameter update becomes:
@@
where $v_{t}^{(i)}$ tracks historical gradients for component $\mathcal{M}_{i}$. This adaptive approach demonstrates superior convergence properties compared to RSGD [43]. These techniques significantly reduce computational overhead while maintaining the benefits of hyperbolic geometry.
@@
2.5 Main challenges in building hyperbolic LLMs
Despite the theoretic-al advantages of hyperbolic geometry for modeling hierarchical data, developing practical hyperbolic large language models faces several fundamental challenges that significantly impact their scalability.
(1) Numerical stability and representation capacity
One of the most critical challenges in hyperbolic LLMs stems from the inherent numerical instabilities of hyperbolic representations in floating-point arithmetic. As demonstrated by Mishne et al. [76], the exponential volume growth property of hyperbolic space leads to catastrophic numerical problems in floating-point arithmetic.
Poincaré ball limitations: Under 64-bit floating-point arithmetic, the Poincaré model exhibits severe representation constraints. Points can only be accurately represented within a ball of radius $r_{0}\approx 38$ (where $r_{0}$ is the maximum stable radius under 64-bit floating-point precision), as the boundary proximity (when $\|\mathbf{x}\|\to 1$) causes numerical instability in the conformal factor $\lambda_{x}=\frac{2}{1-\|\mathbf{x}\|^{2}}$. This leads to gradient vanishing problems where the Riemannian gradient magnitude scales as $\mathcal{O}(\delta^{2})$ for points at distance $\delta$ from the boundary. As demonstrated by Mishne et al. [76], the maximum representable distance corresponds to $|\mathcal{K}|\approx 16$ in the equation $1-10^{-|\mathcal{K}|}=2^{-53}$, severely limiting the model’s capacity to represent distant hierarchical relationships.
Lorentz model constraints: The Lorentz model’s closed-form geodesic computation partially mitigates numerical issues by avoiding division by near-zero denominators in distance calculations:
(12) $d_{\mathcal{L}}(\mathbf{x},\mathbf{y})=\text{arccosh}(-\langle\mathbf{x},\mathbf{y}\rangle_{\mathcal{L}}),$
where $\langle\cdot,\cdot\rangle_{\mathcal{L}}$ denotes the Minkowski inner product. However, the Lorentz model still faces representation capacity limitations within the radius $r_{0}/2\approx 19$.
Sala et al. [92] rigorously characterize a fundamental precision–depth trade‐off for hyperbolic embeddings. They prove that embedding a simple chain graph of length $\ell$ with distortion at most $1+\varepsilon$ requires either: (i) scaling the effective hyperbolic radius linearly in $\ell$; or (ii) allocating $\Theta(\ell/\varepsilon)$ bits of floating-point mantissa precision. Thus, the need to balance numerical precision against the depth (chain length) of the hierarchy is an intrinsic property of hyperbolic geometry.
(2) Computational overhead and scalability
Hyperbolic LLMs face significant computational challenges that limit their practical deployment, particularly for large-scale applications.
Exponential and logarithmic map overhead: The frequent use of exponential and logarithmic mappings between hyperbolic space and tangent space introduces substantial computational overhead [15, 119]. Each transformation requires expensive operations including arcosh, sinh, and cosh functions, leading to training times that are slower than standard Euclidean models. For hybrid architectures, this overhead is amplified by the need for repeated mappings in attention mechanisms, feedforward layers, and normalization operations between Euclidean tangent space and hyperbolic embeddings.
Quadratic complexity in attention: Traditional hyperbolic attention mechanisms suffer from quadratic time complexity $\mathcal{O}(N^{2})$ with respect to the number of input tokens. This scalability bottleneck becomes particularly problematic for long-sequence modeling and large-scale graph data processing.
Memory and precision requirements: Hyperbolic operations often require higher precision arithmetic to maintain numerical stability, leading to increased memory consumption and reduced computational efficiency on hardware accelerators. Multi-component floating-point representations have been proposed as solutions [125], but these approaches introduce additional implementation complexity and may not be efficiently supported on all hardware platforms. To be precise, BigFloats required for computations proposed by Yu et al. [125] are not supported on ML accelerator hardware like GPUs: computations using BigFloats on CPUs will be slow and hence not realistic for many meaningful tasks. Secondly, BigFloats are currently not supported in most deep learning frameworks, even on CPUs.
(3) Training and optimization challenges
Training hyperbolic LLMs presents unique training and optimization challenges that require careful tuning of operations in hyperbolic spaces compared to the training operations happening in standard Euclidean models.
Limited hyperbolic operations: Maintaining points on the hyperbolic manifold throughout training is a critical challenge. Standard optimization steps can easily violate manifold constraints, requiring projection operations or constraint-preserving update schemes. These projections introduce additional computational overhead and can accumulate errors over many training iterations, particularly in deep networks with multiple hyperbolic layers. Many standard neural network operations lack well-defined hyperbolic counterparts. Operations such as batch normalization, dropout, concatenation, and various activation functions require careful geometric reformulation to preserve hyperbolic properties [120, 16]. The absence of these operations has historically limited hyperbolic models to simple architectures, hindering their application to complex tasks that benefit from deep, sophisticated networks.
Curvature selection and adaptation: Choosing appropriate curvature values for different layers and components remains a significant challenge. Fixed curvature models may not optimally represent data with varying hierarchical depths, while learnable curvature approaches introduce additional hyperparameters and optimization complexity [45]. Recent mixture-of-curvature approaches show promise but add substantial architectural complexity.
These challenges collectively represent significant barriers to the widespread adoption of hyperbolic LLMs. However, ongoing research continues to address these limitations through improved numerical methods, more efficient architectures, and better optimization techniques. Understanding these challenges is crucial for researchers seeking to develop robust and scalable hyperbolic language models that can realize the theoretical benefits of hyperbolic geometry in practical applications. Building upon these mathematical foundations, we now present a comprehensive taxonomy of hyperbolic LLM architectures and examine how these geometric principles have been integrated into large language model systems.
@@ 3 Hyperbolic LLMs
A major challenge in developing large language models (LLMs) in hyperbolic space lies in generalizing Euclidean operations to hyperbolic manifolds. Unlike simple element-wise operations (e.g., addition, multiplication, etc.) or fundamental neural network components such as linear transformations, attention mechanisms, and normalization, these operations and components require careful geometric reformulation in hyperbolic space. Even operations like convolution and pooling must be redesigned to respect the curved geometry of hyperbolic space, as standard Euclidean formulations do not preserve the geometric properties essential for hierarchical representation. The fundamental issue stems from the fact that hyperbolic manifolds do not naturally support the linear algebraic operations that underpin modern neural network architectures. This necessitates the development of specialized geometric operations and optimization techniques that respect the curved geometry of hyperbolic space while maintaining computational efficiency. Table 1 presents our taxonomy of these models across four primary categories: hybrid hyperbolic-Euclidean models, hyperbolic fine-tuned models, fully hyperbolic models, and hyperbolic state-space models. Table 2 provides a comprehensive overview of the tasks, dataset types, and specific datasets used to evaluate these HypLLM models across different application domains.
Table 1: Taxonomy of hyperbolic large language models (HypLLMs), organized by architecture type, backbone, model, and geometry.

### v2

@@ 2.2 Exponential and logarithmic mapping
Given a hyperbolic manifold $\mathcal{M}$ (either the Poincaré ball or Lorentz model in $\mathbb{H}^{n}$) and a point $x\in\mathcal{M}$, the tangent space $\mathcal{T}_{x}\mathcal{M}$ corresponds to $\mathbb{R}^{n}$. The exponential map $\exp_{x}:\mathcal{T}_{x}\mathcal{M}\to\mathcal{M}$ maps a tangent vector at $x$ to the point on the manifold $\mathcal{M}$ reached by following the geodesic starting at $x$. The logarithmic map $\log_{x}:\mathcal{M}\to\mathcal{T}_{x}\mathcal{M}$ maps a manifold point to the tangent space at $x$ along the geodesic connecting it to $x$. These bidirectional maps connect the hyperbolic manifold $\mathcal{M}$ and its tangent space $\mathcal{T}_{x}\mathcal{M}$, as shown in Figure 3.
(1) For the Poincaré Ball Model ($\mathcal{M}=\mathcal{B}_{r}^{n}$), the exponential map of vector $v\in\mathcal{T}_{x}\mathcal{M}$ to the manifold $\mathcal{M}$ at any point $x\in\mathcal{M}$ is given by,
(4) $\text{exp}_{x}(v)=x\oplus_{r}\left(\tanh\left(\frac{\lambda_{x}\ \|v\|}{2r}\right)\frac{rv}{\|v\|}\right),$
where $\lambda_{x}$ is the conformal factor and $\oplus_{r}$ denotes Möbius addition. The logarithmic map of $y\in\mathcal{M}$ to the tangent space $\mathcal{T}_{x}\mathcal{M}$ is given by,
(5) $\text{log}_{x}(y)=\frac{2r}{\lambda_{x}}\text{artanh}\left(\frac{\|{-}x\oplus_{r}y\|}{r}\right)\frac{{-}x\oplus_{r}y}{\|{-}x\oplus_{r}y\|}.$
Figure 3: Illustration of exponential and logarithmic mappings between Manifold $\mathcal{M}$ (Poincaré ball or Lorentz model) and the tangent space $\mathcal{T}_{x}\mathcal{M}$ (Euclidean). The exponential map $\exp_{x}(v)$ projects a tangent vector $v\in\mathcal{T}_{x}\mathcal{M}$ onto the manifold, while the logarithmic map $\log_{x}(y)$ projects a manifold point $y\in\mathcal{M}$ back to the tangent space.
(2) For the Lorentz model ($\mathcal{M}=\mathcal{L}^{n+1}_{r}$), the exponential and logarithmic maps at any point $x\in\mathcal{M}$ are defined as follows [14],
(6) $\text{exp}_{x}(v)=\cosh\left(\frac{\sqrt{\langle v,v\rangle_{\mathcal{L}}}}{r}\right)x+r\sinh\left(\frac{\sqrt{\langle v,v\rangle_{\mathcal{L}}}}{r}\right)\frac{v}{\sqrt{\langle v,v\rangle_{\mathcal{L}}}},$
(7) $\text{log}_{x}(y)=d_{\mathcal{L}}(x,y)\frac{y+(r^{-2}\langle x,y\rangle_{\mathcal{L}})x}{\|y+(r^{-2}\langle x,y\rangle_{\mathcal{L}})x\|_{\mathcal{L}}},$
where $d_{\mathcal{L}}(x,y)$ is the hyperbolic distance defined in Equation 3. These mappings preserve radial distances and enable a hybrid workflow: (i) hyperbolic points are mapped to the tangent space via $\text{log}_{x}$, (ii) standard Euclidean operations are performed, and (iii) results are mapped back to hyperbolic space via $\text{exp}_{x}$. This bidirectional mapping between the hyperbolic manifold $\mathcal{M}$ and its tangent space $\mathcal{T}_{x}\mathcal{M}$ enables hybrid models to combine Euclidean computational efficiency with hyperbolic geometric fidelity and hierarchical expressiveness.
To enable direct computation in hyperbolic space, Ungar [108, 107] introduced Möbius gyrovector algebra, enabling vector space operations in the Poincaré ball [40] (see Appendix A). These operations preserve the hyperbolic structure while enabling neural network computations such as linear transformations and activations.
@@ 2.3 Optimization techniques in Riemannian manifold
Training models in hyperbolic space requires specialized optimization techniques, since Euclidean gradient descent is incompatible with curved manifolds. Gradients are computed in the tangent space and then mapped back to the manifold, allowing intrinsic parameter updates [10]. A key insight is to perform gradient updates intrinsically on the manifold via Riemannian optimization [70].
Riemannian Stochastic Gradient Descent (RSGD): RSGD generalizes SGD to Riemannian manifolds by using exponential mappings to update parameters while maintaining manifold constraints [10]:
@@
Riemannian Stochastic Variance Reduced Gradient (RSVRG): Building on this foundation, RSVRG extends the Stochastic Variance Reduced Gradient (SVRG) method to non-Euclidean settings to accelerate convergence. It integrates geometric principles with variance reduction, achieving linear convergence for geodesically convex functions and $\mathcal{O}(1/T)$ convergence for non-convex objectives [131, 70].
Riemannian AdaGrad: Adaptive gradient methods have been extended to Riemannian manifolds [7]. For a product manifold $\mathcal{M}=\mathcal{M}_{1}\times\cdots\times\mathcal{M}_{n}$, the parameter update becomes
@@
where $v_{t}^{(i)}$ tracks historical gradients for component $\mathcal{M}_{i}$, and $\epsilon$ is a small positive constant added for numerical stability to prevent division by zero. This adaptive approach demonstrates superior convergence properties compared to RSGD [7]. These techniques significantly reduce computational overhead while maintaining the benefits of hyperbolic geometry.
@@ 3 Hyperbolic LLMs
Our taxonomy organizes the Hyperbolic LLMs into four categories (Table 1): hybrid hyperbolic-Euclidean models, hyperbolic fine-tuned models, fully hyperbolic models, and hyperbolic state-space models. A summary of HypLLM models, including their tasks, data types, and benchmark datasets across various application domains, is provided in Appendix Table C.1.
Table 1: Taxonomy of hyperbolic large language models (HypLLMs), organized by architecture type, backbone, model name, and geometry.